OpenStudy (anonymous):

Calculate the indefinite integrals, using the substitution method: Anyone know the answer please?

OpenStudy (anonymous):

A) int((x+3)^4, x) b) int(1/(3*x+2), x) c) int(sen(4*x+3), x)

hartnn (hartnn):

what have you already tried for the A) ?

hartnn (hartnn):

did you try u = x+3 ?

OpenStudy (anonymous):

I am not able to deal, I need to deliver this activity today, if anyone can help me, I am grateful :)

hartnn (hartnn):

if u = x+3 what will be du = ... ?

OpenStudy (anonymous):

@fabiomartins are you familiar with steps to solve an integral using substitution?

OpenStudy (anonymous):

could make the resolution step by step, I understand?

OpenStudy (anonymous):